Carol’s Dress Shop produces high quality formal dresses. In January 2019 they produced 17,000 dresses. For the month of January, the following standard and actual cost data are available. The normal monthly capacity of the company is 30,000 direct labor hours. All material purchased in January was used in January production.

Standard per Dress

Actual

Direct materials

5.0 yards @ $8.00 per yard

$660,000 for 80,000 yards

Direct labor

1.5 hours @ $15.00 per hour

$384,000 for 24,000 hours

Overhead

  1. hours @ $5.50 per hour

(fixed $3.40; variable $2.10)

$110,000 fixed overhead

$52,000 variable overhead

Overhead is applied on the basis of direct labor hours. At normal capacity, budgeted fixed overhead costs are $102,000 per month and budgeted variable overhead costs are $63,000 per month.

Required

  1. Calculate the direct materials price variance for January. Label the variance as favorable or unfavorable.
  2. Calculate the direct materials quantity variance for January. Label the variance as favorable or unfavorable.
  3. Calculate the direct labor rate variance for January. Label the variance as favorable or unfavorable.
  4. Calculate the direct labor efficiency variance for January. Label the variance as favorable or unfavorable.
  5. Calculate the variable overhead spending variance for January. Label the variance as favorable or unfavorable.
  6. Calculate the variable overhead efficiency variance for January. Label the variance as favorable or unfavorable.
  7. Calculate the fixed overhead spending variance for January. Label the variance as favorable or unfavorable.
  8. Calculate the fixed overhead production volume variance for January. Label the variance as favorable or unfavorable.
  9. Which of the variances should be investigated if management considers a variance of more than 5% from standard to be significant?
  10. Provide a discussion of the tradeoffs that are most likely to exist between the direct material and direct labor variances.
Answer #1

(1): Direct materials price variance = (actual quantity * actual price) – (actual quantity * standard price)

Actual quantity = 80,000 yards and standard price = $8 per yard.

So, direct materials price variance = 660,000 – (80,000*8)

= 20,000 unfavorable

(2): Direct materials quantity variance = (standard quantity – actual quantity)*standard price

Now standard quantity = 17000*5 = 85,000 yards and actual quantity = 80,000 yards.

Thus direct materials quantity variance = (85000-80000)*$8

= 40,000 favorable

(3): Direct labor rate variance = (actual hours * actual rate) – (actual hours*standard rate)

= 384,000 – (24000*$15)

= 24,000 unfavorable

(4): Direct labor efficiency variance = (standard direct labor hours – actual direct labor hours)*standard rate

= (1.5*17000 – 24000) * $15

= 22,500 favorable

(5): Variable overhead spending variance = (standard variable overhead rate – actual variable overhead rate) * actual units of allocation base

= (2.10 – (52000/24000)) * 24000

= (2.10- 2.16667)*24000

= 1600 unfavorable

(6): Variable overhead efficiency variance = (standard direct labor hours – actual direct labor hours) * standard variable overhead rate

= (1.5*17000-24000)*2.10

= 3,150 favorable

(7): Fixed overhead spending variance = actual fixed overheads – budgeted fixed overheads

= 110,000 – 102,000

= 8,000 unfavorable

(8): Fixed overhead production volume variance = applied fixed overhead – budgeted fixed overhead

Applied fixed overhead = actual output*fixed overhead absorption rate

Fixed overhead absorption rate = 102,000/(30,000/1.5) = 5.10 per dress

Thus fixed overhead production volume variance = (17000*5.10) – 102,000

= 15,300 favorable